# Recognition through Reasoning: Reinforcing Image Geo-localization with Large Vision-Language Models

### v1

@@ 3.3 Model Fine-tuning: GRPO-based Reinforcement Learning
Let $\pi_{\theta}$ denote the current policy parameterized by $\theta$, and let =$\mathcal{B}=\{(\bm{x}_{i},\{\bm{a}_{i}^{(j)}\}_{j=1}^{k})\}$ represent a batch of input prompts $\bm{x}_{i}$ each paired with $k$ candidate completions $\bm{a}_{i}^{(j)}$ sampled from the policy. Each completion $\bm{a}_{i}^{(j)}$ is scored by a composite reward function:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}\in[0,1]$ are weights controlling the importance of the three reward components: locatability ($R_{\text{loc}}$), visual grounding consistency ($R_{\text{vis}}$), and geo-localization accuracy ($R_{\text{geo}}$).
@@
$A_{i}^{(j)}=r_{i}^{(j)}-\frac{1}{k}\sum_{l=1}^{k}r_{i}^{(l)},$ (5)
@@ 4 Experiments
We conduct qualitative and quantitative experiments to evaluate the effectiveness of our curated dataset MP16-Reason and the GRPO-based training strategy used in GLOBE. Specifically, we examine whether MP16-Reason enables better geo-reasoning (i.e., the ability to infer geographic locations through interpretable and visually grounded reasoning) compared to conventional image-only datasets (which lack reasoning supervision) and street-view datasets (which offer limited visual diversity), and whether GRPO training yields stronger reasoning performance than supervised fine-tuning. We also compare GLOBE with both open- and closed-source LVLMs.
@@ 4.1 Experimental Setup
Datasets. The curated dataset MP16-Reason is divided into two subsets: MP16-Reason-Train with 33k samples and MP16-Reason-Test with 12k samples, respectively. MP16-Reason-Train is used to train GLOBE, while MP16-Reason-Test is used to evaluate all baseline methods. To ensure a comprehensive comparison, we additionally evaluate all models on the public geo-localization benchmark IM2GPS3K vo2017revisiting .
@@
Evaluation Metrics. We adopt different evaluation metrics for MP16-Reason-Test and the public geo-localization benchmark to account for differences in annotation format and prediction targets. In MP16-Reason-Test, we evaluate model performance using three metrics: (1) city-level accuracy, (2) country-level accuracy. On the public benchmark, we follow previous work vivanco2023geoclip ; haas2024pigeon ; jia2024g3 ; li2024georeasoner and report the percentage of predictions whose geographic distance to the ground-truth coordinate falls within fixed thresholds (25km, 200km, and 750km). Since our model outputs discrete place names (e.g., country or city), we use external tools to convert these names into their corresponding geographic center coordinates for evaluation.
@@
Implementation details. We implement GLOBE based on the publicly available Qwen2.5-VL-7B Qwen2.5vl , a large vision-language model with strong multimodal understanding capabilities. Instead of using task-specific supervised fine-tuning as a cold start, we directly fine-tune the model using reinforcement learning based on the GRPO framework described in Section 3.3. All training is conducted using MP16-Reason-Train. Further details are provided in Appendix A.1.
@@
Table 1: Localization accuracy on MP16-Reason-Test and IM2GPS3K vo2017revisiting . Underlined results indicate test–train overlap; † denotes models that are not publicly available. Best and second-best results are in bold and blue, respectively.
Method Dataset Size MP16-Reason Test IM2GPS3K vo2017revisiting  (@ km)
City Country City Region Country
Acc. Acc. 25km 200km 750km
ISNs muller2018 MP-16 4M 47.33 82.06 28.00 36.60 49.70
GeoCLIP vivanco2023geoclip MP-16 4M 52.52 84.07 34.47 50.65 69.67
Translocator pramanick2022 MP-16 4M - - 31.10 46.70 58.90
PIGEOTTO haas2024pigeon MP-16 4M - - 36.70 53.80 72.40
G3 (LLaVA) jia2024g3 MP-16 4M - - 35.87 49.42 66.93
G3 (GPT4V) jia2024g3 MP-16 4M - - 40.94 55.56 71.24
Qwen2.5-VL-7B Qwen2.5vl - - 37.33 78.38 32.53 43.14 58.96
Gemma3-27B team2025gemma - - 37.12 72.47 33.37 46.61 63.63
InternVL3-78B zhu2025internvl3 - - 37.31 78.86 35.05 47.32 64.03
Qwen2.5-VL-72B Qwen2.5vl - - 44.45 82.97 35.77 48.35 64.96
GPT-4.1 gpt4-1 - - 53.78 89.87 46.85 60.36 74.41
GeoReasoner li2024georeasoner GSV 133K 32.39 67.30 26.94 36.63 52.27
GaGA dou2024gaga MG-Geo 5M - - 33.00 48.00 67.10
GLOBE (Ours) MP16-Reason 33K 53.16 88.54 40.18 56.19 71.45
4.2 Performance Comparison
We evaluate the effects of the curated dataset and training strategy through quantitative metrics and qualitative analysis, as detailed in the following subsections.
4.2.1 MP16-Reason: Annotation Benefit and Scene Robustness
We validate the contribution of the MP16-Reason dataset through two perspectives. First, we examine whether reasoning-augmented annotations provide tangible benefits over image-only supervision. Second, we compare MP16-Reason to existing geo-reasoning datasets that are predominantly street-view-centric, assessing generalization performance across diverse visual scenes.
Annotation Benefit. To isolate the contribution of our reasoning-based annotations, we ablate the reward components used in GRPO training. Specifically, we compare a model trained only with the geo-localization accuracy reward (VGC Reward), which is derived from the original annotations, as shown in Table 2-row4. As reported in Table 1 and Table 2, GLOBE trained on MP16-Reason significantly outperforms all variants trained on MP-16 (Table 1-row1, Table 1-row2, and Table 1-row3). While PIGEOTTO haas2024pigeon outperforms GLOBE at the country level, its performance lags significantly at the city and region levels (Table 1-row4). It is worth noting that G3 (GPT-4V) jia2024g3 leverages GPT-4V to perform retrieval-augmented generation (RAG) for geo-localization, incorporating external knowledge through a prompt ensembling strategy. Despite this advantage, our method achieves comparable performance at the city level and even slightly outperforms G3 (GPT-4V) at the region and country levels, as shown in Table 1. This highlights the importance of the additional reasoning-based annotations, such as locatability judgments and visual grounding signals, which were distilled from multiple vision-language models and used to supervise finer-grained aspects of geo-reasoning.
Scene Robustness. To assess the generalization ability of geo-reasoning models beyond street-view scenarios, we compare GLOBE with counterpart LVLMs trained on street-view-oriented reasoning datasets (Table 1-row12 and Table 1-row13). While models trained on such datasets perform well on street-view imagery, their generalization degrades considerably in more diverse scenarios (see Table 1). In contrast, MP16-Reason, which includes both street-view and non-street-view images, achieves consistently strong results across a wider range of visual contexts, better reflecting the complexity of real-world geo-localization tasks.
Figure 5: Reasoning comparison of four different models (GPT-4.1 gpt4-1 , GLOBE, Qwen2.5-VL-7B Qwen2.5vl with SFT, and InternVL3-78B zhu2025internvl3 ) on the same input image. Reliable visual clues identified by the models are marked in text.
4.2.2 [Uncaptioned image] GLOBE: Model Performance and Training Strategy Analysis
To further evaluate the effectiveness of our proposed model GLOBE, we conduct comprehensive experiments from two perspectives: external baseline comparison and internal strategy ablation. Externally, we benchmark GLOBE against vision-only models without reasoning supervision, general-purpose LVLMs, and task-specific LVLMs trained on street-view reasoning datasets. Internally, we benchmark SFT against GRPO-based reinforcement learning, and perform ablation studies within the GRPO framework to assess the effect of different reward configurations. These comparisons examine the performance of GLOBE and the effects of its key design choices.
External Baseline Comparison. We compare GLOBE against three categories of representative baselines. The first group includes traditional approaches, which rely on visual feature matching without supporting reasoning capabilities. The second group consists of open-source LVLMs (e.g., Qwen2.5-VL Qwen2.5vl , InternVL3 zhu2025internvl3 , Gemma3 team2025gemma , and GPT-4.1 gpt4-1 ) trained on general-purpose data, which are expected to exhibit broader and more generalizable reasoning capabilities. The third group includes task-specific LVLMs trained on geo-reasoning datasets oriented toward street-view imagery. As shown in Table 1, GPT-4.1 gpt4-1 demonstrates outstanding performance on this task, but it remains closed-source. Aside from GPT-4.1 gpt4-1 , the proposed GLOBE achieves strong accuracy on both the MP16-Reason-Test and the public benchmark IM2GPS3K vo2017revisiting , while also producing more coherent and interpretable reasoning trajectories (see Table 1 and Figure 5). Notably, GLOBE achieves this performance with just 33K samples from MP16-Reason, highlighting the efficiency of reasoning-aware supervision.
Table 2: Ablation study on training methods and reward modeling configurations in MP16-Reason-Test set using Qwen2.5-VL-7B-Instruct model.
Models CoT SFT GRPO Metrics
Loc Reward VGC Reward GA Reward City Country
Acc. Acc.
Qwen2.5-VL-7B 37.21 78.28
Qwen2.5-VL-7B 46.38 84.92
GLOBE w/o Loc&GA 48.80 83.65
GLOBE w/o Loc&VGC 50.27 87.31
GLOBE w/o Loc 50.47 87.15
GLOBE w/o VGC 51.68 87.62
GLOBE w/o GA 49.11 86.53
GLOBE 53.16 88.54
Internal Strategy Ablation. We further investigate the impact of different training paradigms and reward configurations on the performance of GLOBE. Specifically, we compare full-parameter SFT, and our proposed GRPO-based reinforcement learning approach (See Table 2-row2 and Table 2-row8). For GRPO, we conduct a reward ablation by evaluating different combinations of the three reward components: Locatability (Loc) Reward, Visual Grounding Consistency (VGC) Reward, and Geo-localization Accuracy (GA) Reward. It demonstrates that GRPO with the complete reward set achieves the highest overall performance (Table 2-row8). Removing either Loc, VGC or GA Reward results in noticeable drops, underscoring the importance of reasoning-aware supervision beyond location-level correctness (Table 2-row5, Table 2-row6, Table 2-row7). In addition, GRPO demonstrates clear advantages over SFT for the image geo-localization task in LVLMs, offering improved consistency and grounding by directly optimizing the relative quality of generated outputs (Table 2-row2 vs. Table 2-row8).

### v2

@@ 3.3 Model Fine-tuning: GRPO-based Reinforcement Learning
Let $\pi_{\theta}$ denote the current policy parameterized by $\theta$, and let $\mathcal{B}=\{(\bm{x}_{i},\{\bm{a}_{i}^{(j)}\}_{j=1}^{k})\}$ represent a batch of input prompts $\bm{x}_{i}$ each paired with $k$ candidate completions $\bm{a}_{i}^{(j)}$ sampled from the policy. Each completion $\bm{a}_{i}^{(j)}$ is scored by a composite reward function:
@@
where $\lambda_{1},\lambda_{2},\lambda_{3}\in[0,1]$ are weights controlling the importance of the three reward components: localizability ($R_{\text{loc}}$), visual grounding consistency ($R_{\text{vis}}$), and geo-localization accuracy ($R_{\text{geo}}$).
@@
$A_{i}^{(j)}=\frac{r_{i}^{(j)}-\mu_{i}}{\sigma_{i}},\quad\mu_{i}=\frac{1}{k}\sum_{l=1}^{k}r_{i}^{(l)},\quad\sigma_{i}=\sqrt{\frac{1}{k}\sum_{l=1}^{k}\big(r_{i}^{(l)}-\mu_{i}\big)^{2}},$ (5)
@@ 4 Experiments
We conduct both qualitative and quantitative experiments, including ablation studies, to evaluate the effectiveness of our curated dataset MP16-Reason and the GRPO-based training strategy employed in GLOBE. Specifically, we examine whether MP16-Reason enables better geo-reasoning (i.e., the ability to infer geographic locations through interpretable and visually grounded reasoning) compared to conventional image-only datasets (which lack reasoning supervision) and street-view datasets (which offer limited visual diversity). We further assess whether GRPO training provides stronger reasoning capability than supervised fine-tuning, and compare GLOBE against both open- and closed-source LVLMs.
@@ 4.1 Experimental Setup
Datasets. The curated dataset MP16-Reason is divided into two subsets: MP16-Reason-Train with 33k samples and MP16-Reason-Test with 12k samples, respectively. MP16-Reason-Train is used to train GLOBE, while MP16-Reason-Test is used to evaluate all baseline methods. The detailed statistics of these subsets, including sample size, geographic coverage, and scene distribution, are summarized in Table 1. Notably, MP16-Reason-Test was deliberately constructed to cover a broader geographic range (e.g., more countries and cities), including locations not present in the training set, which allows evaluation of how well the model generalizes in geo-reasoning beyond the training distribution. To ensure a comprehensive comparison, we additionally evaluate all models on the public geo-localization benchmark IM2GPS3K vo2017revisiting and OSV-5M astruc2024openstreetview .
Table 1: Statistics of the proposed MP16-Reason.
Dataset #Samples #Country #City #Indoor Scene #Natural Scene #Urban Scene
MP16-Reason-Train 33721 134 1944 5393 2077 26251
MP16-Reason-Test 12000 145 3012 2096 1092 8812
# denotes the number of instances.
@@
Evaluation Metrics. We follow previous work vivanco2023geoclip ; haas2024pigeon ; jia2024g3 ; li2024georeasoner and report the percentage of predictions whose geographic distance to the ground-truth coordinate falls within fixed thresholds (1km, 25km, 200km, 750km, and 2500km). Since our model outputs discrete place names (e.g., country or city), we concatenate the predicted city and country into a single string and query Microsoft Azure Maps  [^1] https://portal.azure.com/, which returns the corresponding representative GPS coordinate (e.g., the geographic center of the region) for evaluation.
@@
Implementation details. For data curation, we deployed Qwen2.5-VL-72B and InternVL3-78B using 8 $\times$ H20 GPUs under the VLLM framework, while GeoCLIP was run separately on a single H20 GPU. These models were used to perform inference over the original MP16 dataset. We then built GLOBE on top of Qwen2.5-VL-7B Qwen2.5vl , a publicly available LVLM with strong multimodal understanding capabilities. Instead of using task-specific supervised fine-tuning as a cold start, we directly fine-tune the model using reinforcement learning based on the GRPO framework described in Section 3.3. In GRPO training, the 7B model was trained on 8 $\times$ H20 GPUs with a batch size of 16, yielding a throughput of approximately 0.44 examples per second. Further implementation details are provided in Appendix A.1.
4.2 Experimental Results
To assess the impact of the curated dataset and the proposed training strategy, we conduct both external baseline comparisons and internal ablation studies, as detailed in the following subsections.
4.2.1 Baseline Comparison
We evaluate the geo-localization performance of GLOBE on both MP16-Reason-Test and the public benchmark IM2GPS3K vo2017revisiting (see Table 4.2.1). For clarity, we organize the baselines into three categories: (I) Image-only supervision, which relies purely on visual features and coordinate labels without reasoning signals; (II) Open- and closed-source LVLMs, including general-purpose LVLMs trained on diverse multimodal data; and (III) Task-specific reasoning supervision, which refers to models trained on geo-localization datasets with reasoning-oriented annotations, often dominated by street-view imagery. We further assess generalization on the street-view dataset OSV-5M astruc2024openstreetview (mini-3K), with detailed results provided in Appendix A.2.2. In addition, we examine performance under different scene conditions to demonstrate the robustness of GLOBE, as reported in Appendix A.2.1.
Table 2: Geo-localization performance comparison on MP16-Reason-Test and IM2GPS3K vo2017revisiting .
Method Dataset, Size MP16-Reason-Test (% @ km) IM2GPS3K vo2017revisiting  (% @ km)
Street City Region Country Continent Street City Region Country Continent
1km 25km 200km 750km 2500km 1km 25km 200km 750km 2500km
[…]
ISNs muller2018 MP-16, 4M 26.24 47.38 55.88 68.48 80.92 10.50 28.00 36.60 49.70 66.00
GeoCLIP vivanco2023geoclip MP-16, 4M 29.28 52.52 66.85 84.07 93.33 14.11 34.47 50.65 69.67 83.82
[…]
Hybrid astruc2024openstreetview OSV-5M, 5M 0.97 16.53 28.72 50.31 71.47 0.83 13.28 25.33 43.84 65.63
RFM-YFCC dufour2025around Flickr, 48M 11.72 46.64 60.46 77.97 91.96 5.41 29.70 44.71 61.83 79.55
[…]
Qwen2.5-VL-7B Qwen2.5vl - 15.42 52.72 62.86 75.11 83.47 8.58 32.53 43.11 58.93 72.37
InternVL3-8B zhu2025internvl3 - 12.01 44.17 55.66 75.36 86.98 6.44 25.69 34.57 49.38 61.66
Gemma3-27B team2025gemma - 16.03 55.63 68.07 82.59 91.29 8.48 33.37 46.61 63.63 79.95
InternVL3-78B zhu2025internvl3 - 14.72 52.46 65.25 81.73 91.17 8.93 35.05 47.32 64.03 78.64
Qwen2.5-VL-72B Qwen2.5vl - 17.52 59.30 71.01 84.06 91.65 9.11 35.77 48.35 64.96 78.88
[…]
GeoReasoner-7B li2024georeasoner GSV, 133K 10.06 40.44 50.91 68.01 79.68 7.67 26.94 36.63 52.27 65.39
[…]
GLOBE-7B (Ours) MP16-Reason, 33K 17.99 62.85 73.83 86.68 92.52 9.84 40.18 56.19 71.45 82.38
@@ 4.2.1 Baseline Comparison
† denotes models that are not publicly available. Underlined results indicate test–train overlap. Best open- and closed-source results are in blue and bold, respectively.
Image-only supervision. Compared with models trained solely on large-scale image-only supervision (e.g., MP-16 with over 4M samples), GLOBE achieves comparable or even superior accuracy using only 33K samples from MP16-Reason. This efficiency gain stems from reasoning-driven supervision, which provides explicit localizability judgments and visual grounding signals beyond raw coordinates. These findings suggest that reasoning annotations can substantially compensate for data scale, enabling more data-efficient geo-localization.
Open- and closed-source LVLMs. GLOBE achieves stronger results than open-source LVLMs, outperforming much larger models such as Qwen2.5-VL-72B Qwen2.5vl and InternVL3-78B zhu2025internvl3 . Notably, GLOBE, built on the Qwen2.5-VL-7B Qwen2.5vl backbone, surpasses Qwen2.5-VL-72B Qwen2.5vl , the larger model originally used to generate the distilled annotations. This outcome highlights the effectiveness of our distillation and GRPO-based training framework in extracting and refining knowledge rather than merely replicating model outputs. In addition, qualitative comparisons of reasoning trajectories are provided in Appendix A.2.5, further illustrating the interpretability advantages of GLOBE. Compared with closed-source industrial systems such as Doubao1.5-VL guo2025seed1 and GPT-4.1 gpt4-1 , GLOBE remains behind. This gap is expected, as the training data scale and settings of these systems are not publicly disclosed. We aim to advance open, reproducible, and data-efficient LVLM training to support sustainable progress.
Task-specific reasoning supervision. Relative to models trained on task-specific reasoning datasets dominated by street-view imagery, GLOBE demonstrates stronger robustness. By incorporating different types of scenes during the construction of MP16-Reason, our approach achieves superior generalization, particularly when evaluated under diverse scene conditions (see Appendix A.2.1). Under comparable 7B backbones, GLOBE consistently outperforms counterparts (Qwen2.5-VL-7B Qwen2.5vl and GeoReasoner-7B li2024georeasoner ), confirming the necessity of scene diversity for real-world geo-localization.
We further evaluate on OSV-5M astruc2024openstreetview (mini-3K), a street-view dataset outside the training domain of MP16-Reason (see Appendix A.2.2). Despite this domain shift, GLOBE surpasses open-source methods such as ISNs muller2018 and GeoCLIP vivanco2023geoclip , which are trained on data distributions similar to MP16-Reason, as well as counterparts such as Qwen2.5-VL-7B Qwen2.5vl and InternVL3-8B zhu2025internvl3 . These results demonstrate that reasoning-driven supervision enhances in-domain performance while enabling superior generalization to unseen domains. Representative failure cases are discussed in Appendix A.2.3, providing qualitative insights into the model’s limitations. Beyond accuracy and generalization, we also provide an efficiency comparison in Appendix A.2.4.
Table 3: Ablation on reward components with Qwen2.5-VL-7B Qwen2.5vl backbone.
Model CoT SFT GRPO MP16-Reason-Test (% @ km)
Loc VGC GA Street City Region Country Continent
Reward Reward Reward 1km 25km 200km 750km 2500km
Qwen2.5-VL-7B Qwen2.5vl 14.37 51.11 61.29 73.67 82.46
Qwen2.5-VL-7B Qwen2.5vl 15.42 52.72 62.86 75.11 83.47
Qwen2.5-VL-7B Qwen2.5vl 16.38 56.76 70.21 83.82 90.75
GLOBE w/o Loc&GA 17.01 59.36 71.77 84.44 91.76
GLOBE w/o Loc&VGC 17.24 59.24 71.93 84.69 91.54
GLOBE w/o Loc 17.50 59.58 71.23 84.06 91.23
GLOBE w/o VGC 17.52 59.83 72.22 84.72 91.12
GLOBE w/o GA 17.44 59.53 71.41 84.33 91.18
GLOBE 17.99 62.85 73.83 86.68 92.52
Best results are in blue.
4.2.2 Ablation Study
To better understand the contributions of our design choices, we conduct ablation studies along three dimensions: (I) the reward components used in GRPO training; (II) the backbone models on which our method is applied; and (III) the distillation datasets employed for supervision. These experiments allow us to disentangle the effects of supervision signals, model capacity, and data quality, thereby providing a more comprehensive view of the strengths of GLOBE and MP16-Reason.
Reward components. Table 3 presents ablation results for GRPO training under different reward configurations, including Localizability (Loc), Visual Grounding Consistency (VGC), and Geo-localization Accuracy (GA). Using all three rewards yields the highest overall performance (row 9). Removing any single component (rows 6-8) causes noticeable drops, highlighting the importance of reasoning-driven supervision beyond coordinate accuracy alone. Moreover, GRPO outperforms SFT (row 9 vs. row 3), delivering stronger consistency and grounding by leveraging reward signals to guide output quality. Even with partial reward combinations, GRPO still surpasses SFT, demonstrating the clear advantage of reinforcement learning with reasoning-driven supervision. In addition to the choice of reward components, the weighting hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ in the GRPO objective also play a role in balancing supervision. A detailed discussion of their design rationale and experimental evaluation is provided in Appendix A.2.6, while an analysis of reward trajectories throughout the training process is presented in Appendix A.2.7.
Table 4: Ablation on backbone architectures.
Backbone Training Strategy MP16-Reason-Test (% @ km)
Street City Region Country Continent
1km 25km 200km 750km 2500km
InternVL3-8B zhu2025internvl3 Baseline 12.01 44.17 55.66 75.36 86.98
SFT 12.41 44.68 56.37 75.20 86.32
GRPO 17.47 60.09 72.41 85.02 91.92
Qwen2.5-VL-7B Qwen2.5vl Baseline 15.42 52.72 62.86 75.11 83.47
SFT 16.38 56.76 70.21 83.82 90.75
GRPO 17.99 62.85 73.83 86.68 92.52
Backbone models. Across both Qwen2.5-VL-7B Qwen2.5vl and InternVL3-8B zhu2025internvl3 , GRPO consistently yields clear improvements over SFT at all geographical levels (see Table 4), confirming the robustness of the training framework. Nevertheless, the absolute performance is influenced by the backbone itself, with Qwen2.5-VL-7B Qwen2.5vl achieving higher post-GRPO accuracy than InternVL3-8B zhu2025internvl3 . These results indicate that GRPO provides stable relative gains across architectures, while the final performance ceiling is determined by backbone capacity and pretraining quality.
Table 5: Ablation on data curation with Qwen2.5-VL-7B Qwen2.5vl backbone.
Curation Setting Training Strategy MP16-Reason-Test (% @ km)
Street City Region Country Continent
1km 25km 200km 750km 2500km
Baseline - 15.42 52.72 62.86 75.11 83.47
Random sampling SFT 15.23 52.00 64.56 78.17 85.23
GRPO 17.26 59.22 71.80 84.73 91.26
Single-source validation SFT 15.22 52.47 65.09 78.79 86.15
GRPO 17.37 59.45 71.88 84.74 91.24
Full multi-source validation SFT 16.38 56.76 70.21 83.82 90.75
GRPO 17.99 62.85 73.83 86.68 92.52
Distillation datasets. To evaluate the contribution of our validation steps in data curation (MP16-Reason-Train), we compare models trained on different curation settings, all standardized to 33K samples but obtained through different filtering strategies, such as random sampling or validation by only a single LVLM (InternVL3-78B zhu2025internvl3 ). As shown in Table 5, these ablated settings lead to noticeable performance drops compared with the full MP16-Reason-Train, confirming the importance of comprehensive validation in constructing a high-quality dataset. In addition, GRPO consistently outperforms SFT across all settings, further highlighting the effectiveness of reinforcement learning in leveraging reasoning-driven supervision.